OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). 5x + y = 5 3x + 2y = 3

OpenStudy (anonymous):

y= (3-3x)/2 therefore 5x + 3/2 -(3x/2) = 5 therefore 3.5x = 3.5 x=1 thus y = 0 (1,0)